Find the slope of any line perpendicular to the line through points (0,5) and (-3,-4)
m=-4-5 over -3-0
m=-9 over -3
m=-3 which is the slope

Answer by venugopalramana(3286)

Find the slope of any line perpendicular to the line through points (0,5) and (-3,-4)
m=-4-5 over -3-0
m=-9 over -3
m=-3 which is the slope ...CORRECT..BUT THIS IS THE SLOPE OF THE GIVEN LINE
ITS PERPENDICULAR WILL HAVE SLOPE = -1/M=-1/-3=1/3